OpenStudy (anonymous):

cosӨ-cosӨsin^2Ө= a) sec^2Ө b) tan^2Ө c)cos^3Ө d)sinӨ

OpenStudy (anonymous):

c)cos^3Ө

OpenStudy (anonymous):

cosӨ(1-sin^2Ө) ---> cosӨ(sin^2Ө+cos^2Ө-sin^2Ө) ---> cosӨcos^2Ө =)cos^3Ө

OpenStudy (anonymous):

Thanks again! You're a beast haha

OpenStudy (anonymous):

\[\cos \theta(1-\sin ^{2}\theta)=\cos \theta.\cos ^{2}\theta=\cos ^{3}\theta \]